OpenStudy (anonymous):

What is the value of c in the equation 2c + 5(c − 5) = 10? 0 2 5 10

OpenStudy (anonymous):

well... qhat do you do first?

OpenStudy (anonymous):

5 * 5?

OpenStudy (anonymous):

wrwrite the whole equation... like where do u put the \(5\times 5\)

OpenStudy (anonymous):

and u are half right by the way

OpenStudy (anonymous):

2c + 25c = 10?

OpenStudy (anonymous):

you distribute the 5(c-5)

OpenStudy (anonymous):

that would be 2c+5(c*5)

OpenStudy (anonymous):

if it's 25c

OpenStudy (bee_see):

Think PEMDAS

OpenStudy (anonymous):

and so if you distribute that... you get 5c-25 right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

that gives you 2c+5c-25=10

OpenStudy (anonymous):

when what do you do?

OpenStudy (anonymous):

then*

OpenStudy (anonymous):

2c+5c = 7c

OpenStudy (anonymous):

7c - 25 = 10

OpenStudy (anonymous):

you simplify the 2c+5c right? you get 8c-25=10

OpenStudy (anonymous):

7c*

OpenStudy (anonymous):

THE ANSWER IS 5 RIGHT

OpenStudy (anonymous):

I GOT IT YESS

OpenStudy (anonymous):

THANKS FOR ALL THE HELP

OpenStudy (anonymous):

then you move the 25 to the other side.. .or add 25 to both sides

OpenStudy (anonymous):

ok :D np

OpenStudy (anonymous):

ITS 5?